40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 4 20:13:39 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors_1_Feb_2019_2230_PST.pdfETMY_UL.pdfETMY_sensors_timeDomain.pdf
    Reply  Thu Feb 7 19:34:18 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors.pdfpostDiag.pdf
       Reply  Fri Feb 8 20:48:52 2019, gautam, Update, LSC, IFO recovery IMG_7307.JPG
          Reply  Mon Feb 11 15:41:49 2019, gautam, Update, LSC, TRY 60 Hz solved - but clipping persists 
             Reply  Mon Feb 11 16:38:34 2019, gautam, Update, LSC, ETMY OL calibration updated OL_calib_ETMY_PERROR.pdfOL_calib_ETMY_YERROR.pdf
             Reply  Fri Feb 15 18:42:57 2019, gautam, Update, LSC, TRY 60 Hz solved  IMG_7323.JPG
Message ID: 14445     Entry time: Fri Feb 8 20:48:52 2019     In reply to: 14441     Reply to this: 14446
Author: gautam 
Type: Update 
Category: LSC 
Subject: IFO recovery 

Several housekeeping tasks were carried out today in preparation for the Y-arm loss measurement.

  1. The mess around the OMC rack was cleared a bit. The vertex laptop paola now lives there, instead of on the ITMY optical table.
  2. Centering of beam on AS photodiodes on AS table (starting from the first optic in this path at the exit point from the vacuum), adjusted AS camera to bring the spot roughly to the center.
  3. POX/POY locking was restored, GTRY/GTRX levels are healthy. TRY was centered on the Thorlabs PD by triggering the LSC lock on AS110.
  4. Oplevs on all four TMs and BS were centered for post-vent alignment.
  5. ETMY OL transfer function was checked since we have swapped the HeNe during the vent, 4.5 Hz UGF for both DoFs and ~30 deg phase margin. The calibration of the error point to urad needs to be double checked.
  6. There are some huge 60 Hz harmonics in the TRY signal - hunting down the source of this. The one thing I can think of that was changed is that we plugged the c1auxey eurocrate into the ethernet powerstrip, I wonder if this created some kind of ground loop.
    • I checked the signal from the PD with a battery powered scope, no evidence of any 60 Hz in the time domain or scope FFT (Attachment #1, FFT in red and time domain signal in green can be seen).
    • Restored the power of c1auxey eurocrate to its original socket in the back of 1Y4 - harmonics still present --> points to the problem being in the whitening board / ADC electronics?
    • The harmonics only seem to show up when TRY > ~0.5
    • Some elog hunting revealed that this signal is being digitized through a modified D990399. So somehow the signal pollution is happening inside this board? Because from the output of this board, the signal is going straight into the ADC.
    • To confirm, I will temporarily hijack another ADC channel and look at the spectrum. There is apparently some kind of daughter board (D040060), but how 60 Hz is coupling at this stage is unknown to me.
  7. The ASS system for both arms still isn't working properly, to be investigated. The dirty TRY signal probably isn't helping the situation.
Attachment 1: IMG_7307.JPG  1.467 MB  Uploaded Sat Feb 9 21:23:34 2019  | Hide | Hide all
IMG_7307.JPG
ELOG V3.1.3-